a father is now four times as old as his son. If the sum of their ages ten years ago was sixty. find thrie present ages
Answer by rapaljer(4671)

Let x = son's age
4x = father's age

x-10 = son's age 10 years ago
4x-10= father's age 10 years ago

x-10 + 4x-10=60
5x-20=60
5x=80
x=16 son's age now
4x=64 father's age now (or was that his grandfather??)

Check: 10 years ago, son was 6 and father was 54, and the sumof their ages is 60.

R^2